sarah123:

I need help solving this C= 5 9 (F−32) The equation above shows how temperature F, measured in degrees Fahrenheit, relates to a temperature C, measured in degrees Celsius. Based on the equation, which of the following must be true? A temperature increase of 1 degree Fahrenheit is equivalent to a temperature increase of 5 9 degree Celsius. A temperature increase of 1 degree Celsius is equivalent to a temperature increase of 1.8 degrees Fahrenheit. A temperature increase of 5 9 degree Fahrenheit is equivalent to a temperature increase of 1 degree Celsius. A) I only B) II only C) III only D) I and II only

sarah123:

@babygirlimarie @BabyGurl2006 @LandonGoma @LILYLUV @Astrid1

sarah123:

I need help on this so i need help

LandonGoma:

If you think of the equation as an equation for a line. y=mx+b where C=5/9(F−32) or C=5/9F−5/9(32)

Astrid1:

If you actually do need help, Do you know how to solve for Celcius to Fahrenheit and Fahrenheit to Celcius? if not Here are some steps: 1.Start with the temperature in Fahrenheit Subtract 30 from this figure (e.g., 100 - 30 = 70). Divide your answer by 2 (e.g., 70 / 2 = 35). (this is gonna be partially off the internet) (I will put my resources below..) Sources:https://reference.yourdictionary.com/resources/what-s-the-easiest-way-to-convert-fahrenheit-to-celsius.html

sarah123:

Ok thanks bro

xxgavinsthunderxx:

bro? lmao

xxgavinsthunderxx:

dats a white *****